2. Suppose that the US and UK are the only countries in the world and that labor is the only productive input.

Labor Requirements for

One unit of X One unit of Y

UK 24 16

US 12 12

a. Which country has absolute advantage in X? in Y? Explain.

b. Which country has comparative advantage in X? in Y? Explain.



1
Expert's answer
2021-04-06T06:55:07-0400

a. Great Britain has an absolute advantage in commodity X (24>12) and Y (16>12).


b. X

UK "\frac{24}{12}=2" - comparative advantage in X


US "\frac {12}{2}=0.5"


Y

UK "\frac{16}{12}=1.3" comparative advantage in Y


US "\frac {12}{16}=0.75"